### v2

# Continual-MEGA: A Large-scale Benchmark for Generalizable Continual Anomaly Detection

Abstract
In this paper, we introduce a new benchmark for continual learning in anomaly detection, aimed at better reflecting real-world deployment scenarios. Our benchmark, Continual-MEGA, includes a large and diverse dataset that significantly expands existing evaluation settings by combining carefully curated existing datasets with our newly proposed dataset, ContinualAD. In addition to standard continual learning with expanded quantity, we propose a novel scenario that measures zero-shot generalization to unseen classes—those not observed during continual adaptation. This setting poses a new problem setting that continual adaptation also enhances zero-shot performance. We also present a unified baseline algorithm that improves robustness in few-shot detection and maintains strong generalization. Through extensive evaluations, we report three key findings: (1) existing methods show substantial room for improvement, particularly in pixel-level defect localization; (2) our proposed method consistently outperforms prior approaches; and (3) the newly introduced ContinualAD dataset enhances the performance of strong anomaly detection models. We release the benchmark and code in https://github.com/Continual-Mega/Continual-Mega.
1 Introduction
Anomaly detection (AD) cao2024adaclip ; defard2021padim ; deng2022anomaly ; gudovskiy2022cflow ; huang2024adapting ; jeong2023winclip ; liu2023simplenet ; roth2022towards ; strater2024generalad ; yao2024hierarchical ; yao2023one ; you2022unified ; zhang2023destseg ; zhou2023anomalyclip plays a critical role in quality control, ensuring precise identification of defects during production. It is widely used in sectors such as manufacturing, where automatic anomaly detection can significantly improve operational efficiency and product safety. Due to the complexity of real-world environments, anomaly detection models are required to recognize a diverse range of defects. To address the issue, many scenarios with public datasets bergmann2019mvtec ; zou2022spot ; mishra2021vt ; wang2024real ; lehr2024viaduct ; lehr2024ad3 have been proposed. Conventional deep approaches bergmann2022beyond ; cohen2020sub ; defard2021padim ; li2021cutpaste ; ristea2022self ; roth2022towards ; zavrtanik2021draem ; zou2022spot suppose unsupervised or per-class anomaly detections. Following the advancement of CLIP radford2021learning and its initial application to anomaly detection jeong2023winclip , unified anomaly detection frameworks you2022unified ; he2024learning ; yao2024hierarchical ; strater2024generalad have emerged, enabling a single model to handle diverse evaluation scenarios. These approaches include continuous learning and adaption li2022towards ; pang2025context ; liu2024unsupervised ; tang2024incremental ; jin2024oner ; meng2024moead ; mcintosh2024unsupervised as well as zero-, few-shot anomaly detection jeong2023winclip ; zhou2023anomalyclip ; li2024promptad ; zhu2024toward ; deng2023anovl ; chen2024clip ; chen2023zero ; tamura2023random ; gu2024filo ; gui2024few ; qu2024vcp .
From a dataset perspective, the inherent difficulty of collecting large numbers of samples, particularly defective ones, makes anomaly detection (AD) more challenging than general vision tasks. Consequently, widely used evaluation datasets bergmann2019mvtec ; zou2022spot are notably limited in both size and variability. This limitation would have motivated recent research to explore continual, zero-shot, and few-shot learning settings as strategies to overcome the scarcity of data. Reflecting the need, we suggest the need of new evaluation benchmarks with expanded data quantity, achieved by integrating diverse public datasets and curating additional samples to increase both the volume and variety of data.
In this paper, we introduce a novel and comprehensive evaluation benchmark, Continual-MEGA, that evaluates the continual and zero-shot capabilities of anomaly detection models. Our benchmark includes a large-scale evaluation dataset that integrates widely used public datasets bergmann2019mvtec ; lehr2024viaduct ; lehr2024ad3 ; mishra2021vt ; wang2024real with a newly curated dataset, ContinualAD. The Continual-MEGA dataset supports two primary evaluation scenarios: (1) a standard continual learning setup, and (2) an extended setup for evaluating generalization performance after the continual learning phase, often required in applications. To further validate the effectiveness of our curated dataset, ContinualAD, we also introduce an additional scenario where the ContinualAD dataset is excluded from training.
The extensive evaluation conducted on the proposed Continual-MEGA benchmark, we test the representative anomaly detection methods cao2024adaclip ; huang2024adapting ; liu2024unsupervised ; liu2023simplenet ; qu2024vcp ; strater2024generalad ; tang2024incremental ; tao2024kernel ; yao2024hierarchical ; zhang2024mediclip ; zhou2023anomalyclip and clearly demonstrate that there is still substantial room for improvement in the AD domain in perspective of continual adaptation and generalizability. Additionally, we propose a new baseline AD method for the Continual-MEGA benchmark that enables parameter-efficient and continuous adaptation of pre-trained CLIP. Our approach applies mixture-of-expert (MoE) style multi-layered-perceptron (MLP) adaptor utilization, combined with anomaly feature synthesis and fine-tuning of prompt embeddings, reporting the state-of-the-art performance in the proposed Continual-MEGA benchmark, expected to act as a new baseline method for future research.
Our contributions are summarized as follows:
  • We introduce Continual-MEGA, a novel large-scale continual learning benchmark, featuring detailed evaluation scenarios. The benchmark is constructed by integrating existing public datasets with our newly curated dataset, ContinualAD, which notably expands the overall data volume and diversity.
  • From extensive evaluations on the proposed Continual-MEGA benchmark, we demonstrate that there is still enough room for improvement in the performance of AD algorithms.
  • We propose a new anomaly detection method called Anomaly Detection across Continual Tasks (ADCT) that combines MoE-style adaptor modules, anomaly feature synthesis, and prompt-based feature tuning with CLIP. Our method achieves state-of-the-art performance on the Continual-MEGA benchmark and would serve as a strong baseline for future research.
Figure 1: Example visualizations of sample images from various public anomaly detection datasets and the proposed ContinualAD dataset. Green boxes indicate normal images, while red boxes represent anomaly images.
2 Continual-MEGA Benchmark
Datasets Composition.
We compose a new benchmark to evaluate continual learning for anomaly detection in large-scale real-world settings, including various public datasets including MVTec-AD bergmann2019mvtec , VisA zou2022spot , Real-IAD wang2024real , VIADUCT lehr2024ad3 , BTAD mishra2021vt , and MPDD jezek2021deep , and also with the newly proposed ConitnualAD dataset, which consists of diverse images collected from real-world objects. Figure 1 shows example images from the seven datasets included in the proposed Continual-MEGA benchmark, demonstrating the diversity and complexity of anomaly types across domains. The ConitnualAD dataset consists of a total of 30 classes, comprising 14,655 normal images and 15,827 anomaly images, significantly larger than widely used MVTec-AD and VisA datasets, as illustrated in Figure 2. To evaluate the continual learning performance of comparative methods, we design two experimental scenarios. The model is initially pre-trained on either 85 or 58 classes, followed by continual learning with 60 novel classes introduced incrementally.
Dataset Acquisition.
The proposed ContinualAD dataset images were collected using 10 devices: Galaxy S21+, iPhone 12 Pro Max, iPhone 13, iPhone 15 Pro Max, iPhone XS, iPad Air 4, iPad Pro 11-inch 2nd generation, iPad Pro 12.9-inch 4th generation, iPhone 12 mini, and ZFLIP 3. This diversity in devices and capturing conditions enables evaluation under a wide range of real-world anomaly scenarios and environmental variations.
Various Scenarios for Continual Learning.
[…]
Scenario 1 extends conventional continual learning settings by combining the MVTec-AD bergmann2019mvtec and VisA zou2022spot datasets, widely used in anomaly detection. We pretrain the model on all 858585 Base classes and sequentially introduce 555, 101010, and 303030 New classes over 121212, 666, and 222 iterations, respectively. Scenario 2 is designed to evaluate zero-shot generalization following continual adaptation. In this setting, both MVTec-AD and VisA are excluded from the continual learning process, they are neither part of the Base nor New classes. Instead, they are held out solely for assessing the model’s zero-shot performance, serving as a novel protocol to evaluate cross-domain generalization. Additionally, Scenario 3 further analyzes the generalization capability of the proposed ContinualAD dataset by removing the target dataset from the Base classes and New classes stream. Specifically, the model is continually adapted with 303030 New classes from other datasets, while zero-shot generalization is evaluated on the excluded dataset, following the setup of Scenario 2.
Figure 3 demonstrates the detailed statistics for each of the two main scenarios. For both cases, we can see that there exists an imbalance for each classes. Considering that we measure the amount of forgetting class-wise, we can suppose that it would be advantageous for fitting smaller classes from previous datasets bergmann2019mvtec ; zou2022spot . We investigate the supposition by comparing the results from Scenario 1 and 2 in the Experiment section.
Figure 2: Illustration of statistics of various datasets. Each graph (from left to right) shows the number of classes, number of images, and pixel value variance for public datasets, as well as our proposed ContinualAD and Continual-MEGA.
(a) Class distribution of Scenario 1.
(b) Class distribution of Scenario 2.
Figure 3: Class distributions of different scenarios. Each colored region denotes a Base, and each divided (New) tasks. The volume in the orange line in the histogram represents the number of anomaly samples over the entire sample volume represented by gray. Detailed class configuration for each task will be presented in the Supplementary material.
Metrics and Implementation Details.
For all quantitative evaluations, we adopt two metrics proposed in tang2024incremental to evaluate continual learning performance in anomaly detection: average accuracy (ACC) and forgetting measure (FM). The ACC metrics are computed on the basis of the image-level area under the ROC curve (AUROC) and the pixel-level average precision (AP) liu2024unsupervised , providing a comprehensive view of both the classification accuracy and the model’s resilience to forgetting over time. For the FM measure, we evaluate the accuracy drop for each ACC after the adaptation. All the methods are trained and evaluated in equivalent training settings with 50 epochs for baseline method training, and 20 epochs for continual adaptation for fair comparison, following concerns raised in cha2024hyperparameters regarding the fair comparison by the hyper-parameter settings of continual learning evaluation. For the proposed baseline method, a single set of hyperparameters was derived through lightweight tuning on the base classes of Scenario 1 and then consistently applied across all remaining scenarios without further adjustment. This avoids per-scenario tuning and better reflects realistic continual learning settings, where fine-grained tuning for each incoming task is typically impractical.
3 Proposed Baseline Method
(a) Training scenario of $Base$ and Task $N$.
(b) Inference process. $B$ and $N$ denote the adapters corresponding to the base classes and the task-specific classes $\{1,2,3,...,N\}$, respectively. To inference, we use $A^{\text{avg}}$, which is the average of the adapter weights trained on the base classes and the $N$ task adapters.
Figure 4: Overview architecture.
3.1 Mixture-of-Expert of Adapters
To solve the Continual-MEGA benchmark, we propose a new baseline AD method, using CLIP radford2021learning text and visual encoders. The diagrams in Figures 4(a) and 4(b) represent the overall framework of the method. We use a set of four-adapter $A=\{A_{1},\dots A_{4}\}$ for each block of layers $S_{1}$ to $S_{4}$ of the CLIP visual encoder. For each categories including the set of (Base) classes $C_{b}$ and the set of $n$’th task $C_{n}$ classes, where the entire task number is $n=1,\dots N$, we separately train the adapter set denoted as $A_{n}=\{A_{1,B},\dots,A_{4,B}\}$. The features $F_{1}$ to $F_{4}$ from the adapters are added to form features to represent the anomaly map, and we apply the proposed loss function for training all the adapters. Specifically, we use the Text encoder converted features for preset normal and abnormal text prompts to supplement AD training. The text features $F_{\text{text}}\in\mathbb{R}^{2\times d}$ for normal and anomaly are obtained from a text encoder through the text prompt $p$. The detailed settings for the text prompt are available in the appendix A.3.
In the inference stage, we accumulate all the pre-trained adapters $A_{n}=\{A_{1,B},\dots,A_{4,B}\}$ and $A_{b}$ for each task and baseline class set, by the average adapters $A^{\text{avg}}=\{A^{\text{avg}}_{1},...,A^{\text{avg}}_{4}\}$, as follows:
$A^{\text{avg}}_{l}=\frac{1}{N+1}(\sum^{N}_{n}{A_{l,n}}+A_{l,b}).\end{aligned}$ (1)
where the number $l=\{1,\dots 4\}$ denotes the ordering of each of four blocks. In implementation, each adaptation layer $A_{l}(\cdot)$ consists of two linear layers as:
$A_{l}(F_{l})=W_{l,2}(W_{l,1}F_{l}^{T}),$ (2)
where $F_{l}\in\mathbb{R}^{G\times d}$ represents the visual features extracted from the visual encoder stage $S_{l}$, whereas $G$ and $d$ present the number of grids and feature dimension of each grid, respectively. $W_{l,1}$ and $W_{l,2}$ denote the learnable parameters of linear transformations. To preserve the original knowledge from the pre-trained CLIP model, the input $\tilde{F}_{l}$ to $S_{l+1}$ is designed by a residual connection (gao2024clip, ) as
$\tilde{F}_{l}=\alpha F_{l}+(1-\alpha)A_{l}(F_{l}),$ (3)
where $\alpha$ represents a residual ratio between the original features and the adapted features. $\alpha$ is set to 0.9 in our work. We use the CLIP with ViT-L/14 dosovitskiy2020image architecture, which consists of 24 sublayers divided into four layers, where each layer contains six sublayers. The size of input images was set to 336. The adaptation layers for anomaly feature generation were applied to layers 1, 2, 3, and 4.
3.2 Synthetic Feature Generation
Specifically, we apply random noise to enable the adaptation layers $A_{l}$ to learn a diverse range of anomalies. In training, we use task-wise zdators $A_{n}$ and in the inference phase, we use the accumulated adators $A_{avg}$. The synthetic anomaly features ($F_{l}^{1}$) are generated by
$F_{l}^{1}=A_{l}(F_{l}+\gamma),$ (4)
where $\gamma\in\mathbb{R}^{G\times d}$ is a random noise. The adapted normal features ($F_{l}^{0}$) are generated via the adaptation layers as
$F_{l}^{0}=A_{l}(F_{l}).$ (5)
The adapted normal features $F_{l}^{0}$ and synthetic anomaly features $F_{l}^{1}$ are booth used to generate anomaly score maps through cosine similarity operations along with the text features.
3.3 Model Training
The proposed loss function $L(\cdot)$ are implemented to detect pixel-wise anomalies. For each layer $l$, we define pixel-wise losses $L_{\text{no}}$, $L_{\text{an}}$, $L_{\text{syn}}$, each representing the losses for normal, anomalies, and synthetic features, as:
$\mathcal{L}_{\text{no}}=L_{\text{ce}}+L_{d}+L_{f},$ (6)
$\mathcal{L}_{\text{an}}=\mathcal{L}_{d}+L_{f},$ (7)
$\mathcal{L}_{\text{syn}}=L_{\text{ce}}.$ (8)
The losses termed as $L_{\text{ce}}$, $L_{f}$ and $L_{d}$ are the pixel-level cross-entropy, focal (ross2017focal, ) and dice  (milletari2016v, ) loss. All three losses get features $F_{l}$ as input, and get feature mask $M_{l}$ as label. For the mask $M_{l}$, the real-anomaly region and synthetic feature region are masked with ones, and the real-normal region is masked as 0. Finally, the total loss for the layer $l$ is computed as the summation of the previously defined losses in equation (6), (7) and (8), as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{no}}+\mathcal{L}_{\text{an}}+\mathcal{L}_{\text{syn}}.$ (9)
For training, we accumulate all the layerwise loss functions to update adapters.
4 Related Works
Anomaly detection is a specialized task focused on detecting and rejecting unknown samples amodei2016concrete ; hendrycks2021unsolved , often framed as an out-of-distribution (OOD) problem, particularly targeting industrial anomalous data chalapathy2019deep ; pang2021deep ; ruff2021unifying . Specifically, anomaly detection involves both identifying image-level anomalies and localizing defective regions. However, object categories exhibit diverse characteristics, and the detection challenges vary significantly across these categories. Due to the nature, early deep anomaly detection models suppose per-class, or unsupervised anomaly detection bergmann2022beyond ; cohen2020sub ; defard2021padim ; li2021cutpaste ; ristea2022self ; roth2022towards ; zavrtanik2021draem ; zou2022spot .
Toward Unified Anomaly Detection.
Recent advances in large-scale backbone models, such as CLIP radford2021learning , offer a promising solution to the challenge of unified anomaly detection across categories jeong2023winclip ; he2024learning ; yao2024hierarchical ; strater2024generalad . Following the initial approach jeong2023winclip utilizing CLIP, current trends are focused on developing unified anomaly detection models with zero- and few-shot adaptation capabilities across categories.
Zero- and Few-shot Adaptation.
Anomaly detection with zero- and few-shot adaptation jeong2023winclip ; li2024promptad ; chen2023zero ; tamura2023random ; gu2024filo ; gui2024few ; qu2024vcp ; zhou2023anomalyclip ; deng2023anovl ; chen2024clip across various categories reflects real-world scenarios where acquiring a sufficient number of samples for newly incoming categories is often not feasible, and obtaining anomaly samples is even more challenging. In the few-shot adaptation scenario, typically one to five normal or anomaly samples are used to fine-tune the original adaptation models. The zero-shot adaptation scenario assumes no adaptation across categories. Various methods have been proposed to address these challenges, including text prompt utilization jeong2023winclip ; li2024promptad ; zhou2023anomalyclip ; deng2023anovl ; chen2024clip ; tamura2023random ; gu2024filo , visual context prompting qu2024vcp ; deng2023anovl , and anomaly dataset synthesis chen2023zero ; chen2024unified . Notably, most of these approaches leverage text prompt information, with strategies ranging from manually designed templates jeong2023winclip ; deng2023anovl , normal sample-only li2024promptad , learned zhou2023anomalyclip ; deng2023anovl ; gu2024filo , to augmented prompting tamura2023random .
Continual Adaptation.
Another important aspect in recent anomaly detection research trend is continual adaptation li2022towards ; pang2025context ; liu2024unsupervised ; tang2024incremental ; jin2024oner ; meng2024moead ; mcintosh2024unsupervised , which mirrors the scenario where object categories arrive incrementally. In this context, we aim to mitigate catastrophic forgetting while ensuring that adaptation to previous categories improves the model’s performance for future category adaptations. Building on initial efforts li2022towards , several approaches have been proposed, including context-aware feature adaptation pang2025context , learned text prompts liu2024unsupervised , integration of a unified reconstruction-based detection framework tang2024incremental , online replay memory design jin2024oner , parameter-efficient tuning meng2024moead , and unsupervised tuning mcintosh2024unsupervised ; tang2024incremental . The continual evaluation scenario is built on widely-used public datasets such as MVTec-AD bergmann2019mvtec or VisA zou2022spot , but the quantity and diversity of the dataset are limited compared to the continual adaptation scenario bang2021rainbow using ImageNet deng2009imagenet .
Building upon existing studies, the core novelty of our work lies in introducing a new research agenda for continual anomaly detection. First, we propose Continual-MEGA, a large-scale evaluation benchmark designed to measure overall AD performance in more challenging and scalable scenarios. Second, we present a novel AD method that integrates efficient CLIP-based adaptation with anomaly feature synthesis and optimized prompt tuning. Unlike prior approaches such as chen2024unified ; chen2023zero ; meng2024moead , our method demonstrates robust performance across the newly proposed benchmark, highlighting its effectiveness under more demanding evaluation settings.
5 Experiments
Overview.
Tables 1 and 2 present the quantitative results under the proposed evaluation scenarios, comparing various recently proposed anomaly detection methods. In our experiments, we categorize the methods into three groups: (1) approaches that adapt using only normal samples: SimpleNet liu2023simplenet , GeneralAD strater2024generalad , HGAD yao2024hierarchical , and ResAD yao2024resad , (2) vision-language model (VLM)-based methods: MVFA huang2024adapting , VCP-CLIP qu2024vcp , and MediCLIP zhang2024mediclip , and (3) methods specifically designed for continual learning settings: UCAD liu2024unsupervised , and IUF tang2024incremental . Overall, the results indicate a substantial drop in performance across all methods—particularly for pixel-wise anomaly detection—when evaluated under the proposed continual learning settings. This contrasts sharply with the higher performance typically observed in standard benchmarks such as MVTec-AD and VisA, highlighting the increased difficulty and practical relevance of our evaluation protocol.
Evaluation of Continual Learning Capability.
Scenario 1 represents a typical continual learning setup but significantly scales up both the number of classes and data volume compared to prior works on continual adaptation liu2024unsupervised ; tang2024incremental . Notably, vision-language model (VLM)-based methods such as MVFA zhang2024mediclip and MediCLIP zhang2024mediclip achieve the highest performance among all baselines, apart from the proposed method. Specifically, MVFA and our proposed method achieved comparable performance to each other. The overall results strongly suggest that the existing methods, including continual learning approaches for anomaly detection, struggle to handle large-scale continual evaluation settings.
Table 1: Experimental results on Scenario 1. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. While all methods were trained with the same number of epochs for fair comparison, the $\text{IUF}^{\ast}$ method requires substantially longer training due to its methodology. Therefore, we trained the $Base$ classes for 500 epochs and the $New$ classes for 100 epochs in the $\text{IUF}^{\ast}$ setting. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
  Type Method 85-5 (12 tasks) 85-10 (6 tasks) 85-30 (2 tasks)
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$)
Only-normal SimpleNet 56.5/4.0/30.3 7.1/2.7/4.9 56.4/4.3/30.4 6.2/2.4/4.3 58.2/4.5/31.4 2.4/1.8/2.1
GeneralAD 49.3/1.5/25.4 5.5/1.2/3.4 50.2/1.4/25.8 3.2/1.5/2.4 48.9/1.1/25.0 5.8/1.2/3.5
HGAD 54.1/5.2/29.7 1.5/0.4/1.0 53.3/5.3/29.3 2.1/0.3/1.2 52.7/5.3/29.0 4.8/0.0/2.4
ResAD 73.1/13.9/43.5 1.3/0.4/0.8 71.9/12.7/42.3 1.0/0.3/0.6 70.3/10.1/40.2 0.2/1.5/0.8
VLM-based MVFA 75.4/24.4/49.9 4.2/5.6/4.9 76.4/24.3/50.4 4.0/6.8/5.4 75.7/24.8/50.3 6.3/10.3/8.3
VCP-CLIP 44.1/19.3/31.7 2.5/9.0/5.7 61.9/25.6/43.7 4.4/4.1/4.2 44.7/28.9/36.8 4.0/2.4/3.2
MediCLIP 80.5/8.8/44.7 1.4/6.0/3.7 77.9/6.9/42.4 2.2/10.2/6.2 77.7/9.7/43.7 0.4/20.0/10.2
Continual UCAD 67.1/10.8/39.0 0.2/0.0/0.1 64.6/7.8/36.2 0.3/0.03/0.2 57.9/4.4/31.2 1.2/0.0/0.6
IUF 59.8/5.8/32.8 1.3/0.3/0.8 60.1/6.0/33.1 0.1/0.1/0.1 59.8/5.9/32.9 0.5/0.4/0.5
$\text{IUF}^{\ast}$ 61.5/7.4/34.5 0.5/0.3/0.4 61.4/7.6/34.5 0.5/0.1/0.3 63.0/8.8/35.9 0.4/0.3/0.4
Ours 73.8/25.7/49.8 2.0/2.1/2.1 75.8/28.0/51.9 1.3/1.9/1.6 78.9/32.7/55.8 0.8/1.8/1.3
This observation reveals two key insights: (1) several prior methods appear to be tightly fitted to existing benchmarks such as MVTec-AD and VisA, which limits their generalizability to more diverse or challenging settings, and (2) methods with stronger initial (pretrained) performance tend to retain higher accuracy throughout continual adaptation. Regarding the first insight, methods such as MVFA, which demonstrate competitive performance in Scenario 1, exhibit significantly degraded results, particularly in pixel-level AP, when MVTec-AD and VisA datasets are excluded from the Base classes and New classes, as shown in Table 2. In contrast, our proposed method consistently achieves robust performance across all evaluation scenarios. In our proposed setup, detecting the baseline categories is substantially more challenging than in conventional benchmarks, as shown in Table 3. Under this setting, VLM-based methods demonstrate significantly stronger performance compared to methods explicitly designed for continual learning. This discrepancy can be attributed to the limited detection capability of existing continual anomaly detection methods, even at their initial stage, a factor closely tied to the second insight discussed earlier. Consequently, these methods face greater difficulty in adapting to new incoming categories. Although the forgetting measure (FM) of continual learning-based methods appears lower than that of VLM-based methods, this is likely due to their poor initial detection performance rather than effective forgetting mitigation.
Table 2: Experimental results on Scenario 2. To evaluate the zero-shot generalization performance of the methods, we excluded the MVTec-AD and VisA classes from training and used them only for evaluation. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. Zero-shot performance on MVTec-AD and VisA is presented in Figure 5. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
  Type Method 58-5 (12 tasks) 58-10 (6 tasks) 58-30 (2 tasks)
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$)
Only-normal SimpleNet 56.1/4.2/30.2 8.2/1.4/4.8 56.5/3.8/30.2 7.1/2.4/4.8 57.3/3.8/30.6 4.9/1.0/3.0
GeneralAD 49.0/0.8/24.9 6.3/1.7/4.0 51.3/0.9/26.1 3.1/1.2/2.2 47.7/2.1/24.9 5.7/0.0/2.9
HGAD 51.1/4.3/27.7 1.8/0.3/1.1 51.8/4.5/28.2 1.4/0.2/0.8 51.8/4.3/28.1 2.4/0.2/1.3
ResAD 48.8/0.6/24.7 10.7/1.0/5.8 42.7/1.7/22.2 3.0/0.9/1.9 55.6/12.8/34.2 12.0/4.7/8.3
VLM-based MVFA 63.2/4.7/34.0 5.8/5.4/5.6 64.0/4.1/34.1 5.8/2.9/4.4 65.3/5.0/35.2 1.9/2.0/2.0
VCP-CLIP 55.6/18.7/37.1 3.8/6.8/5.3 53.2/19.8/36.5 0.3/3.0/1.7 64.3/22.3/48.3 2.8/3.7/3.3
MediCLIP 79.6/7.3/43.5 3.8/5.6/4.7 76.0/6.0/41.0 4.9/3.7/4.3 77.1/5.9/41.5 2.1/7.0/4.6
Continual UCAD 66.0/7.4/36.7 0.4/0.02/0.2 63.5/6.0/34.8 0.7/0.03/0.4 58.0/3.1/30.6 0.0/0.0/0.0
IUF 57.6/4.2/30.9 1.7/0.5/1.1 58.0/4.3/31.2 0.3/0.2/0.3 58.0/4.3/31.2 -0.7/-0.1/-0.4
$\text{IUF}^{\ast}$ 60.2/6.3/33.3 0.8/0.3/0.6 60.7/6.4/33.6 0.2/0.1/0.2 61.7/7.0/34.4 0.2/0.2/0.2
Ours 71.7/20.7/46.2 2.3/4.1/3.2 72.4/22.2/47.3 2.5/3.8/3.2 76.8/27.5/52.2 1.0/2.6/1.8
Evaluation of Generalizability after Continual Adaptation.
Another notable contribution of our proposed evaluation protocol is the incorporation of zero-shot generalization evaluation following continual adaptation, defined as Scenario 2 in Table 2. This setting reflects practical situations where models must retain or even improve zero-shot capability during continual learning. However, this vital aspect has been largely overlooked in previous anomaly detection benchmarks, which do not adequately capture the demands of real-world deployment. As shown in Table 2 and the accompanying visualizations in Figure 5, only a few methods, most notably MediCLIP, demonstrate decent generalizability after adaptation. We conjecture that meaningful analysis of zero-shot trends is only possible for methods that already exhibit sufficient generalization capability prior to adaptation. Specifically, by comparing results with Scenario 3 in Figure 5, where the proposed ContinualAD dataset is excluded from the Base classes and New classes, we observe that only methods with robust performance in continual and zero-shot generalization settings exhibit performance gains when additional categories are introduced. The full comparison results for the continual learning setting in Scenario 3 are provided in the appendix B.
Summarization.
The overall quantitative results highlight three key takeaways. First, under our expanded evaluation setup, most existing methods, including those explicitly designed for continual learning, exhibit insufficient detection performance, limiting their utility for meaningful analysis. Specifically, the pixel-level AP reported under our proposed evaluation setting shows a substantial performance gap compared to existing benchmarks, with significantly lower scores across most methods. Second, a small subset of methods, primarily VLM-based approaches, demonstrate relatively strong performance and benefit from decent generalization after continual adaptation. From this perspective, Scenario 3 further supports the value of our proposed ContinualAD dataset, showing that it provides meaningful supervision signals that improve detection capability for future incoming categories. Most importantly, the proposed baseline method consistently outperforms existing approaches across all scenarios, establishing a strong benchmark and emphasizing the significance of both the proposed method and the newly introduced continual evaluation protocol.
(a) Scenario 2.
(b) Scenario 3.
Figure 5: Zero-shot performance comparison on MVTec-AD and VisA under Scenario 2 and Scenario 3. The upper two plots present results from Scenario 2, while the remaining plots correspond to Scenario 3. Each point denotes the performance of an AD method, with image-level AUROC on the x-axis and pixel-level AP on the y-axis. Best viewed in wide.
Table 3: Experimental results of base classes across scenarios. We note that the base classes used in Scenario 2 and Scenario 3 differ, as ContinualAD is included among the base classes in Scenario 2 but not in Scenario 3. The best-performing results are highlighted in bold.
  Type Method Scenario 1 (85 classes) Scenario 2 (58 classes) Scenario 3 (58 classes)
Image Pixel Image Pixel Image Pixel
Only-normal SimpleNet 58.8 6.3 61.3 4.5 57.5 4.5
GeneralAD 51.5 2.6 52.6 1.8 54.4 2.7
HGAD 59.5 5.0 56.1 3.2 55.5 2.7
ResAD 73.3 15.5 69.1 7.8 70.7 15.3
VLM-based MVFA 81.7 32.6 65.8 10.4 70.7 21.2
VCP-CLIP 73.8 25.4 61.0 23.1 61.9 22.5
MediCLIP 73.9 4.5 78.1 8.5 75.3 5.9
Continual UCAD 55.8 1.6 58.1 4.7 56.0 3.6
IUF 60.5 7.4 57.4 4.4 58.5 4.2
$\text{IUF}^{\ast}$ 68.3 13.5 65.8 9.5 63.6 9.5
Ours 83.1 39.0 82.0 35.7 77.8 36.5
6 Conclusion
We present Continual-MEGA, a new large-scale benchmark for continual anomaly detection (AD), built by integrating multiple public datasets and curating a novel dataset, ContinualAD, to significantly enhance sample volume and diversity. Comprehensive evaluations on Continual-MEGA reveal that there is still substantial room for improvement in existing AD methods, highlighting both the need for further research in continual AD and the effectiveness of the curated ContinualAD dataset. Also, we introduce a novel baseline method that integrates MoE-style adaptor modules, anomaly feature synthesis, and prompt-based feature tuning with CLIP. Our approach achieves state-of-the-art performance on Continual-MEGA, providing a strong baseline for future work. Limitation: Since our evaluation benchmark still has class imbalance, enhancing the benchmark by improving category balance and expanding data diversity will improve the Continual-MEGA benchmark.
7 Acknowledgment
This work was supported by the Institute of Information &\&& Communications Technology Planning &\&& Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University) and RS-2022-II220124, Development of Artificial Intelligence Technology for Self-Improving Competency-Aware Learning Capabilities].
Appendix A Continual-MEGA Benchmark Configuration Details
A.1 ContinualAD Dataset
To form the Continual-MEGA benchmark, we propose the ContinualAD dataset which has significantly larger quantities and volumes compared to previous datasets. The ContinualAD dataset consists of a total of 30 classes, as listed in Table A. Table D describes the number of normal and anomaly samples for each class in the ContinualAD dataset. Figure A illustrates sample images from the ContinualAD dataset, which includes diverse scenes captured in different background settings. Moreover, the ContinualAD dataset includes a wide range of object instances within the same class, enabling the evaluation of robustness to intra-class variation.
Table A: Class list of ContinualAD dataset.
       Class Names
Energy-bar Apple Kleenex Ruler Toothpaste
Sunglasses Capsule Flash-drive Band-aid Cucumber
Toothbrush Soap Dollar Pencil Fork
Toy Multi-pen Chopsticks Watermelon Egg
Spoon Calculator Eraser Mango Candy
Mouse Glasses-case Notebook Food-container Cup
Consequently, as shown in Figure 2, the image variance of the ContinualAD dataset is significantly higher than that of existing datasets. The variance value is computed by first calculating the pixel-wise variance across images within each class, and then averaging these values across all classes in the dataset. This indicates that prior datasets mainly consist of highly similar images within each class, limiting their ability to evaluate model performance under diverse conditions. In contrast, the higher variance in ContinualAD facilitates more realistic and challenging evaluation scenarios.
Figure A: Example visualization of ContinualAD dataset. For comprehensive benchmarking across diverse environments, the ContinualAD dataset was curated to encompass images featuring a wide range of backgrounds. The red boxes indicate the anomaly regions.
A.2 Training Setting for Continual-MEGA Benchmark
To simulate a low-resource environment where training samples are limited, we adopt a minimal supervision setting in the proposed Continual-MEGA benchmark. Specifically, only 10 normal and 10 anomalous training images are provided per class during both the base and continual learning stages. Table C shows the number of training and test images used for base classes and continual learning classes in each scenario. For continual learning, the classes are partitioned into three task settings, with each task comprising 5, 15, and 30 classes, respectively, to simulate varying levels of incremental difficulty. Following recent trends in AD toward few-/zero-shot and continual learning, we adopt a limited number of training and adaptation samples to better reflect realistic constraints. However, depending on the target AD application, the training setup of our ContinualAD benchmark can be flexibly reconfigured to suit different deployment scenarios. This will be discussed in the Limitation section in more detail.
For the proposed baseline method, hyperparameter tuning was conducted solely on the base classes of Scenario 1 in a lightweight manner. The resulting hyperparameters were uniformly used across all remaining scenarios to ensure fair and consistent evaluation. To evaluate the model performance in a realistic continual learning setting, we avoided scenario-specific hyperparameter tuning on purpose. This design choice aim to reflect practical constraints in real-world deployments, where care tuning for each newly incoming task is often infeasible cha2024hyperparameters .
A.3 Text Prompting Details
We used generalized text prompts to obtain text features that are not specific to any particular domain or class. Table B shows the generalized text prompts used to obtain general text features. It consists of 10 prompts for both normal and anomaly classes, respectively.
Table B: Description of generalized text prompts.
  Prompt No. Normal Prompts Anomaly Prompts
1 This is an example of a normal object This is an example of an anomalous object
2 This is a typical appearance of the object This is not the typical appearance of the object
3 This is what a normal object looks like This is what an anomaly looks like
4 A photo of a normal object A photo of an anomalous object
5 This is not an anomaly This is an example of an abnormal object
6 This is an example of a standard object This is an example of an abnormal object
7 This is the standard appearance of the object This is not the usual appearance of the object
8 This is what a standard object looks like This is what an abnormal object looks like
9 A photo of a standard object A photo of an abnormal object
10 This object meets standard characteristics An abnormality detected in this object
Table C: Number of training and test samples in each scenario.
  Scenario Stage Train Test
#Normal #Anomaly #Normal #Anomaly
Scenario 1 Base 850 850 71,274 44,301
Continual 600 600 49,543 28,140
Scenario 2 Base 580 580 59,121 35,972
Continual 600 600 60,267 34,281
Scenario 3 Base 580 580 69,788 37,130
Continual 300 300 35,245 17,597
Appendix B Continual-MEGA Benchmark Evaluation Details
This section discusses the deeper analysis of the meaning of evaluation results of various AD methods on our Continual-MEGA Benchmark presented in Section 5 of the paper. We note that the scenario 3 exclude ContinualAD datasets both for Base and New classes of the benchmark. Figure B compares the zero-shot performance of various methods on the MVTec-AD and VisA datasets under two conditions: (1) trained only on the Base classes, and (2) after continual adaptation as defined by our proposed Continual-MEGA benchmark. Notably, our proposed baseline model demonstrates improved generalization, benefiting from both a stronger set of Base classes and the continual adaptation of additional classes. Scenario 2 includes our proposed ContinualAD dataset, under which most methods exhibit improved zero-shot performance after continual learning. In contrast, when ContinualAD is excluded in Scenario 3, most existing methods suffer a degradation in zero-shot generalization after continual learning. Among them, our proposed method shows the smallest performance drop, indicating stronger robustness to continual adaptation compared to other approaches. To further investigate this observation, we refer to the quantitative results from Scenarios 2 and 3, presented in Table E and Table F. These tables provide detailed evaluation metrics corresponding to continual adaptation and zero-shot evaluation results. For easier display of the results, we visualize the results for representative methods of the tables in Figure 5.
Table D: Number of normal and anomaly samples per class in the ContinualAD dataset.
  Class #Normal #Anomaly Class #Normal #Anomaly
Energy-bar 329 542 Toy 368 492
Apple 490 502 Multi-pen 494 492
Kleenex 480 519 Chopsticks 488 524
Ruler 277 490 Watermelon 497 506
Toothpaste 513 516 Egg 662 491
Sunglasses 499 572 Spoon 527 517
Capsule 507 493 Calculator 506 500
Flash-drive 522 495 Eraser 458 508
Band-aid 511 491 Mango 456 491
Cucumber 505 507 Candy 517 490
Toothbrush 500 549 Mouse 517 495
Soap 394 787 Glasses-case 501 549
Dollar 391 494 Notebook 354 513
Pencil 518 517 Food-container 520 489
Fork 537 507 Cup 517 488
Table E: Experimental results on Scenario 2. To evaluate the zero-shot generalization performance of the methods, we excluded the MVTec-AD and VisA classes from training and used them only for evaluation. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. Zero-shot performance on MVTec-AD and VisA is presented in Figure 5. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
  Type Method 58-5 (12 tasks) 58-10 (6 tasks) 58-30 (2 tasks) zero-shot (Avg.)
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) MVTec-AD VisA
Only-normal SimpleNet 56.1/4.2/30.2 8.2/1.4/4.8 56.5/3.8/30.2 7.1/2.4/4.8 57.3/3.8/30.6 4.9/1.0/3.0 55.8/9.7/32.8 52.6/0.0/26.3
GeneralAD 49.0/0.8/24.9 6.3/1.7/4.0 51.3/0.9/26.1 3.1/1.2/2.2 47.7/2.1/24.9 5.7/0.0/2.9 53.3/5.8/29.6 49.2/1.4/25.3
HGAD 51.1/4.3/27.7 1.8/0.3/1.1 51.8/4.5/28.2 1.4/0.2/0.8 51.8/4.3/28.1 2.4/0.2/1.3 50.1/16.1/33.1 55.1/2.7/28.9
ResAD 48.8/0.6/24.7 10.7/1.0/5.8 42.7/1.7/22.2 3.0/0.9/1.9 55.6/12.8/34.2 12.0/4.7/8.3 69.7/11.1/40.4 57.8/3.1/30.4
VLM-based MVFA 63.2/4.7/34.0 5.8/5.4/5.6 64.0/4.1/34.1 5.8/2.9/4.4 65.3/5.0/35.2 1.9/2.0/2.0 56.1/5.1/30.6 53.8/2.5/28.2
AnomalyCLIP 52.9/2.0/27.5 4.1/0.9/2.5 51.3/1.9/26.6 1.5/0.6/1.1 51.1/2.2/26.7 2.2/0.2/1.2 57.2/7.0/32.1 51.3/3.6/27.5
VCP-CLIP 55.6/18.7/37.1 3.8/6.8/5.3 53.2/19.8/36.5 0.3/3.0/1.7 64.3/22.3/48.3 2.8/3.7/3.3 62.3/22.7/42.5 61.0/11.2/36.1
MediCLIP 79.6/7.3/43.5 3.8/5.6/4.7 76.0/6.0/41.0 4.9/3.7/4.3 77.1/5.9/41.5 2.1/7.0/4.6 84.2/19.1/51.7 74.1/5.2/39.7
Continual UCAD 66.0/7.4/36.7 0.4/0.02/0.2 63.5/6.0/34.8 0.7/0.03/0.4 58.0/3.1/30.6 0.0/0.0/0.0 61.6/9.4/35.5 54.1/1.9/28.0
IUF 57.6/4.2/30.9 1.7/0.5/1.1 58.0/4.3/31.2 0.3/0.2/0.3 58.0/4.3/31.2 -0.7/-0.1/-0.4 68.0/16.2/42.1 54.7/2.8/28.8
$\text{IUF}^{\ast}$ 60.2/6.3/33.3 0.8/0.3/0.6 60.7/6.4/33.6 0.2/0.1/0.2 61.7/7.0/34.4 0.2/0.2/0.2 67.8/15.4/41.6 58.2/4.9/31.6
Ours 71.7/20.7/46.2 2.3/4.1/3.2 72.4/22.2/47.3 2.5/3.8/3.2 76.8/27.5/52.2 1.0/2.6/1.8 78.4/31.5/55.0 76.9/17.2/47.0
From the results presented, we observe that anomaly detection (AD) performance has following tendencies: (1) Compared to Scenario 3, overall AD performance improves in Scenario 2 across most methods, showing the effectiveness of the ContinualAD dataset. (2) Continual adaptation using the ContinualAD dataset enhances zero-shot generalizability, as observed in VisA and MVTec. Excluding ContinualAD leads to a consistent drop in performance among prior methods. (3) Our proposed baseline achieves strong and consistent results across all scenarios, showing notable improvements in Scenario 2 and maintaining competitive generalizability in Scenario 3 after continual adaptation. Based on these results, we can reasonably conjecture that our ContinualAD dataset provides valuable information for detecting anomalies in both unseen and continually introduced categories under more challenging scenarios than prior setups. The proposed baseline exhibits robust and consistent performance, setting a meaningful benchmark for future AD methods.
(a) MVTec-AD.
(b) VisA.
Figure B: Image-level AUROC and pixel-level AP performance on MVTec-AD (top) and VisA (bottom) datasets. Each point represents the performance of a method before (•) and after ($\times$) continual learning. Arrows indicate the performance change from the model trained only on base classes to the model trained via continual learning. The continual learning results are averaged over three settings, where each task consists of 5, 10, and 30 New classes, respectively.
Table F: Experimental results on Scenario 3. To verify the effectiveness of the ContinualAD dataset, we excluded it from the training process. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP, and average value. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
  Type Method 58-5 (6 tasks) 58-10 (3 tasks) 58-30 (1 tasks) zero-shot (Avg.)
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) MVTec-AD VisA
Only-normal SimpleNet 57.6/5.5/31.6 7.2/3.0/5.1 59.5/7.2/33.4 5.9/2.4/4.2 59.8/6.8/33.3 2.2/0.4/1.3 50.2/7.8/29.0 49.9/0.0/25.0
GeneralAD 50.6/0.7/25.7 3.3/2.0/2.7 51.7/1.0/26.3 5.3/2.6/3.9 51.7/1.4/26.6 3.3/0.9/2.1 52.0/6.3/29.2 51.7/2.5/27.1
HGAD 53.2/3.7/28.5 2.5/0.1/1.3 53.2/3.8/28.5 2.9/0.0/1.4 53.4/3.7/28.6 2.7/0.0/1.4 49.5/15.6/32.6 57.4/2.9/30.2
ResAD 44.6/2.3/23.5 1.3/0.3/0.8 40.5/0.8/20.7 7.9/4.3/6.1 64.2/4.1/34.2 7.8/0.8/4.3 78.0/12.1/45.1 67.6/7.6/37.6
VLM-based MVFA 63.3/6.2/34.8 8.1/10.7/9.4 68.0/11.0/39.5 3.8/10.1/7.0 69.6/16.4/43.0 3.7/5.4/4.6 69.7/9.8/39.8 69.8/5.3/37.6
AnomalyCLIP 51.4/2.6/27.0 5.3/1.7/3.5 53.5/2.7/28.1 1.3/0.4/0.9 54.1/3.1/28.6 -1.1/0.2/-0.4 51.7/6.8/29.3 49.9/2.7/26.3
VCP-CLIP 54.3/21.2/37.8 1.9/2.9/2.4 46.1/18.1/32.1 -0.2/3.3/1.6 61.5/21.0/41.3 2.1/1.2/1.6 58.9/22.5/40.7 58.0/10.6/34.3
MediCLIP 77.3/7.1/42.2 3.6/4.6/4.1 76.0/5.0/40.5 2.0/2.9/2.5 73.2/5.3/39.3 6.3/3.5/4.9 81.8/17.3/49.6 74.7/4.3/39.5
Continual UCAD 65.0/9.6/37.3 0.0/0.0/0.0 59.8/5.8/32.8 0.0/0.0/0.0 55.2/3.4/29.3 0.0/0.0/0.0 59.7/9.0/34.3 53.6/1.7/27.7
IUF 58.1/4.6/31.4 1.2/0.4/0.8 57.6/4.4/31.0 0.1/0.2/0.2 57.6/4.2/30.9 0.3/0.1/0.2 67.8/15.8/41.8 54.9/2.7/28.8
$\text{IUF}^{\ast}$ 59.3/6.3/32.8 0.5/0.3/0.4 59.7/6.5/33.1 0.8/0.1/0.5 60.9/7.4/34.2 0.5/0.4/0.5 64.6/14.5/39.6 57.8/3.5/30.7
Ours 69.5/19.7/44.6 3.2/3.4/3.3 72.7/23.1/47.9 2.4/3.7/3.1 76.8/29.5/53.2 -0.3/2.1/0.9 75.0/28.4/51.7 69.7/13.7/41.7
Appendix C Deeper Discussion of the Limitation and Future Research
Regarding the dataset sample configuration, a primary limitation of the proposed benchmark is class imbalance, as sample sizes vary significantly across datasets. In our continual evaluation setup, models that better fit classes having a smaller number of samples would be beneficial to achieve higher performance. While this setting reflects the class imbalance commonly observed in real-world inspection tasks, balancing sample quantities across classes would improve the reliability of AD performance evaluation in the continual setup.
Regarding the Benchmark training and evaluation configuration, the Continual-MEGA benchmark intentionally adopts limited training and adaptation samples to evaluate the effectiveness of recent few-/zero-shot AD methods under both unseen and continual setups, leveraging a significantly larger evaluation set. While our primary focus is evaluation, we expect higher accuracy with increased training data, particularly for the Base set, making detailed analysis across varying training sizes a key direction for future research.
From a model perspective, our baseline, despite its simplicity, achieves strong performance across diverse scenarios in the Continual-MEGA benchmark. As this work focuses primarily on benchmark construction, deeper analysis through ablations and developing improved AD methods remain essential directions for future research.